• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 00:00:23 on August22,2003

    Entry number 112504

    keyword=Swing shift summary

    16:00-20:00 no beam (Hall C current calibration and accelerator problems)
    15:45-16:30 control access (Sirish inserted the Compton electron detector manually)
    20:30 the aspect ratio is 180x170um
    20:30-20:40 DAQ reloading
    20:40 Start data taking (10um, mult.12C, 2.922 GeV/c)
    20:40 Compton tune was optimized
    20:55 Raster tune: 3.5x3.5mm**2 at 15.2Ax35.A
    21:40-22:03 Setting to 2.813 GeV/c
    22:22-22:46 Setting to 2.704 GeV/c
    23:05 - Start Moller tunning
    23:30 - MCC told us that the Compton lock will not work untill morning
    23:45 - Start Moller measurement